What would be the maximum speed with which a car can round a level curve of radius R = 80.0 m on wet concrete? (The coefficient of static friction between the wet concrete and the tire is 0.60.)

correct answer is: 21.7 m/s

please show your work and explain thoroughly
Answer #1

R = 80m

u = 0.6

From centripetal force = friction force

mV^2/R = umg

=) V = (ugR)^0.5 = (0.6×9.8×80)^0.5 = 21.7 m/s

=) V = 21.7m/s